# DiSCo Meets LLMs: A Unified Approach for Sparse Retrieval and Contextual Distillation in Conversational Search

### v1

@@ Abstract.
Conversational Search (CS) is the task of retrieving relevant documents from a corpus within a conversational context, combining retrieval with conversational context modeling. With the explosion of Large Language Models (LLMs), the CS field has seen major improvements with LLMs rewriting user queries, accounting for conversational context. However, engaging LLMs at inference time harms efficiency. Current methods address this by distilling embeddings from human-rewritten queries to learn the context modeling task. Yet, these approaches predominantly focus on context modeling, and only treat the contrastive component of the retrieval task within a distillation-independent loss term. To address these limitations, we propose a new distillation method, as a relaxation of the previous objective, unifying retrieval and context modeling. We relax the existing training objectives by distilling similarity scores between conversations and documents, rather than relying solely on representation learning. Our proposed distillation objective allows for more freedom in the representation space and leverages the contrastive nature of document relevance. Through experiments on Learned Sparse Retrieval (LSR) across 5 CS datasets, our approach demonstrates substantial improvements in both in-domain and out-of-domain retrieval performance, outperforming state-of-the-art with gains of up to 6 points in recall for out-of-domain datasets. Additionally, through the relaxation of the objective, we propose a multi-teacher distillation, using multiple LLMs as teachers, yielding additional gains, and outperforming the teachers themselves in in-domain experiments. Finally, analysis of the sparsity of the models reveals that our distillation allows for better control over the sparsity of the trained models.
@@ 1. Introduction
Training retrieval models for this task is however challenging, because of the lack of large-scale conversational datasets, making human annotation an essential component of the process. In particular, a lot of effort has been invested first to create conversations with passage relevance judgment (Dalton et al., 2020b; Aliannejadi et al., 2024), and then to rewrite for each user utterance the contextualized version of each query (rewritten query) as the optimal denoised query of each turn (Elgohary et al., 2019). Datasets with rewrites enable to learn the Conversational Query Rewriting (CQR) task by auto-regressive models (e.g., T5 (Raffel et al., 2019)), and then only pass the rewritten queries to retrieval models instead of the full noisy conversations. However, this two-step approach – rewrite and retrieve – is not efficient, and may lead to information loss and error within the rewrite phase that can propagate to the retrieval phase. Hence the need for unified retrieval models that do both tasks together in the representation space (Penha and Hauff, 2020; Mo et al., 2024b).
@@
Approaches such as ConvDR (Yu et al., 2021), coSPLADE (Hai Le et al., 2023) and LeCoRe (Mao et al., 2023c) all combine both conversational context modeling and retrieval tasks within a single retrieval framework, either in a dense or sparse embedding space. They use a distillation objective enforcing the conversation representations to converge to the representations of gold human-rewritten queries. This objective is, however, restrictive and hypothesizes that gold-rewritten query representations are the best and only learning target in the representations space, leaving no freedom for the model to further learn and optimize the representations. Besides, in all these works, the conversational context modeling task has an independent loss function from the ranking loss. We thus propose to fill this gap, with a single distillation loss that unifies both context modeling and ranking tasks, while not limiting the model to learn one single representation. previously used distillation on representations.
@@
In particular, and as described in Figure 1 our method learns to distill the similarity scores between rewritten queries and documents rather than query representations directly. By distilling similarities, we first align with the contrastive nature of the ranking task (Tao and Zhai, 2007; Liu et al., 2021; Wang and Isola, 2020), but also we relax the training objective toward the final goal of retrieval, which is to compute similarities between queries and documents (Weinberger and Saul, 2009; Koch, 2015; Song et al., 2015). This relaxation allows the student to further learn the context modeling task, considering relevant and irrelevant documents from the corpus, rather than mapping all representations to the human representations. This objective also follows the precepts from Hofstätter et al. (2020) on distillation for ad-hoc retrieval, and can benefit from hard negatives within the distillation loss. Both dense and sparse methods could be subject to this relaxation; however, in this work, we focus on learned sparse architectures (Zamani et al., 2018), such as SPLADE (Formal et al., 2021a; Lassance et al., 2024b), as the relaxation would have more degrees of freedom due to the high dimensionality of the representations for sparse retrieval.
@@
Besides, the proposed relaxation also reduces the level of constraints on the teachers, as any method producing similarity scores could be used for the distillation. In the context of CS, this allows the model to learn from multiple LLM rewrites, by fusing similarity scores of several teachers into a single score to be distilled. To the best of our knowledge, our work is the first to distill the knowledge of multiple LLM teachers for query rewriting. This also distinguishes our method from the work from Hofstätter et al. (2020), which distills across architectures, as we distill from the same model but with different inputs: the conversation and the LLM rewrites.
@@
We propose a relaxation of the training objective of CS model, distilling similarities of rewritten queries rather than representations.111We will release our codes and checkpoints upon acceptance.
We extend the method to distill knowledge from multiple teachers, unlocking the potential of LLMs.
We conduct extensive experiments to evaluate the effectiveness of the method on in-domain (QReCC, TopiOCQA) and out-of-domain (TREC CAsT 2020, 2022 and iKAT 2023) datasets.
@@
Our results demonstrate that our Distillation of Sparse Con-versational retrieval, DiSCo, leads to in-domain performance improvements, through our relaxation of the distillation loss. We also see improved generalization capacities with 12% gains on recall and 16% on precision compared to previous zero-shot models (LeCoRe and QRACDR) on CAsT 2020. We demonstrate that relying on multiple LLM teachers for the distillation task brings further performance gains, compared to single teacher model. Finally, thanks to the increased freedom on the representations, we show that we can better control the model sparsity, even for long context, in deeper turns of the conversations.
@@ 2. Related Work
Conversational Search (CS) differentiates itself from ad-hoc search primarily through the nature of the input. Conversational history can grow very long with turns’ dependencies, while search engine queries are typically concise and limited to a few words (Cambazoglu et al., 2021; Lupart et al., 2023). The main challenge in addressing CS is thus modeling the conversational context of the input, to only represent useful information from the past conversation (Wu et al., 2021). Two possibilities exist to learn this noise reduction: first on the token level, with generative models that learn to generate contextualized queries from past conversations (Elgohary et al., 2019), or within the representation spaces (Yu et al., 2021) of retrieval models.
@@
Query rewriting. Learning to resolve the conversational history is difficult, as no or very few conversational search engines are in production today, limiting the availability of user-interaction data. The effort is thus mostly based on human-annotated data. One of the first conversational datasets are QuAC (Choi et al., 2018) and ORConvQA (Qu et al., 2020), where human annotators were tasked to create conversations out of existing documents, assuming a human-machine conversation over the topic of the document. Each paragraph could potentially be an utterance and relevant passages would be inferred by construction. Later CANARD (Elgohary et al., 2019) was released on top of ORConvQA, consisting of human rewrites of each utterance of the conversations, to help learn the context modeling task. This defined one of the main sub-tasks of CS: Conversational Query Rewriting (CQR) (Elgohary et al., 2019; Ye et al., 2023). Following studies learned the CQR task at the token level – generating automatic query rewrites based on the conversation history – on autoregressive language models (e.g., T5 or BART) (Mao et al., 2023a; Wu et al., 2021). Similar methods (Mo et al., 2023) used language models to answer the query, before resolving the context. Today, work is still being done to solve the CQR task, using LLMs in zero- or few-shot fashion. CHIQ (Mo et al., 2024a) proposes to decompose the context modeling task into several simpler sub-tasks for the LLMs – history-enhancement, answer generation, question disambiguation – to gain in interpretability and effectiveness, LLM4CS aggregates several rewrites and answers within the representation space of the retrieval models (Mao et al., 2023b), and MQ4CS focuses on multi-aspect query rewrites (Abbasiantaeb and Aliannejadi, 2024). However, context modeling using LLMs is too computationally costly, making it impossible for production.
@@
Distillation in CS. Lots of research aims at modeling the conversational context via learning to represent the conversation based on gold rewritten queries. Existing methods achieve this goal by taking the representation of the rewritten gold query as a teacher model and learning to distill the representation (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024b). ConvDR (Yu et al., 2021) learns the mapping between full conversation and human rewrites representations by minimizing an Mean Square Error (MSE) loss on the CLS tokens of both, performing well for dense bi-encoders. QRACDR (Mo et al., 2024b), also proposes a similar distillation, with several new MSE terms, between documents and queries to improve the representations and align better with the contrastive nature of the task. We differentiate from these works as we inspire from the contrastive margin distillation (Margin-MSE (Hofstätter et al., 2020)) and relax the query rewriting distillation by learning based on the relevance scores (i.e., the dot product of the query representation and a document), rather than the representation itself. This way, not only do we enable converging towards indefinite possible query representations, but also we take advantage of the contrastive nature of ranking. Furthermore, we are not bound to only one query or one teacher.
@@
Learned sparse representation. As Learned Sparse Retrieval (LSR) gained popularity in ad-hoc retrieval (Lassance et al., 2024b; Lassance and Clinchant, 2023, 2022b), SPLADE architectures were proposed for conversational passage retrieval (Hai Le et al., 2023; Mao et al., 2023c), to benefit their interpretability and robustness properties (Formal et al., 2021b, 2022; Lupart and Clinchant, 2023). In coSPLADE (Hai Le et al., 2023), the authors use an MSE loss between full conversation and human rewrite representations of the sparse bag-of-words representations (of dimension 30k), and show promising performance. In the meantime, LeCoRe (Mao et al., 2023c) also aims to distill human rewrites on sparse representations, through intermediate embedding layers and by filtering a maximum of dimensions, to avoid the MSE on large dimensionality vectors. Distilling sparse representations is however challenging, as it requires controlling which dimensions should be activated, often restricting the potential of the sparse representations. Additionally, using a MSE is problematic since it treats each dimension separately, neglecting the significance of the associated dimension. These issues limit the effectiveness of sparse retrieval in previous work. In our work, by only distilling end scores, we give the model complete freedom to learn which dimensions to activate. We also investigate sparsity further, including methods to control it within CS.
@@ 3.1. Preliminaries
Notation. We consider a set of conversations between a user and a system, each composed of multiple turns. At turn , we have access to past queries and answers, together with the final user utterance . We denote the full conversational context as , and as the gold rewritten query representing the whole conversational context. resolves various complexities such as ellipsis and ambiguities of the conversation and can be generated by either human or LLM.222Prompts used to generate the LLMs rewrites can be found in the Appendix C. We use the [SEP] token as the separator between queries and answers.
Furthermore, similarly to other approaches in CS (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024b) we rely on two already trained encoder models and , for queries and documents representations. These backbone models are trained on a large-scale IR dataset with regular short-form queries extracted from search engine logs. We also follow the assumption made in CS that is already good at representing documents and doesn’t need further fine-tuning, only needs further fine-tuning to adapt to long conversational contexts, noted . This can be done with a contrastive InfoNCE (Oord et al., 2018) loss on the CS datasets (e.g., convANCE, convSPLADE), or with a distillation loss (e.g., ConvDR and other distillation approaches such as DiSCo).
Learned Sparse Retrieval. We rely on the SPLADE architecture (Formal et al., 2021b, a), based on the BERT pretrained transformers (Devlin et al., 2018; Lin et al., 2020). SPLADE makes use of the Masked Language Modeling embedding layer of BERT to create sparse representations, where each dimension corresponds to a token of the vocabulary. , and are the MLM outputs of the model, of dimension 30k (vocabulary dimension). To control the sparsity of the model, the authors propose to use a regularization loss (Paria et al., 2020), The sparsity of the models can be controlled with the two hyperparameters and , as weights for the regularization loss. As we rely on the same SPLADE model, we also included this loss. Further details can be found in the Appendix A.
@@ 3.2. DiSCo
Representation distillation. Existing approaches (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024b) distill query rewrites on the representation space, with a direct constraint on the representation. The goal is to have the representations of the full conversations converging toward the representations of the rewritten queries:
@@
with targeted scores , and the dimension of the representation space. Intuitively, in Figure 2, we illustrate how relaxing the distillation objective can make the learning space easier. While the existing distillation loss functions bound the model to converge to one single embedding (Yu et al., 2021), our loss allows for infinite possible optimum embeddings, as long as they have the same dot product with the target relevant document embedding (i.e., any point in the hyperplane on the figure). Such distillation on scores can be trained with a Kullback Leibler Divergence loss (Kullback and Leibler, 1951) on the distribution of scores:
@@
Hard negatives. Applied to the case of multiple negatives per query. The target space from Equation 2 becomes a subspace of (h-N) dimensions. This still keeps a higher degree of freedom, since the number of negatives is much lower than the dimension of the embedding space ().
@@
Teacher Models. An important component of the distillation relies on the choice of the teacher. The teacher is first used to rewrite the conversation utterance, into , which will be encoded and distilled through the similarities scores . We use multiple LLMs, together with human rewrite (when available) as teachers. This results in several trained DiSCo students, depending on the teacher used. Below we list all the teachers we use:
@@
Furthermore, another property of the proposed distillation is that we can distill from multiple teachers, making the distill from several LLMs or humans possible. We do so by averaging the similarity scores of the teachers, each having a different rewrite, thus a diffe- rent similarity with the documents. Overall, the distilled score is:
@@
Student-Teacher Fusion. Finally, we propose a fusion of the teacher with the student model, combining the ranked lists of both at inference. This is achieved using an average normalized score over the ranked list of both models.

### v2

# by *[inlinelist,1]label=),itemjoin=, ,itemjoin*=, and DiSCo: LLM Knowledge Distillation for Efficient Sparse Retrieval in Conversational Search

@@ Abstract.
Conversational Search (CS) involves retrieving relevant documents from a corpus while considering the conversational context, integrating retrieval with context modeling. Recent advancements in Large Language Models (LLMs) have significantly enhanced CS by enabling query rewriting based on conversational context. However, employing LLMs during inference poses efficiency challenges. Existing solutions mitigate this issue by distilling embeddings derived from human-rewritten queries, focusing primarily on learning the context modeling task. These methods, however, often separate the contrastive retrieval task from the distillation process, treating it as an independent loss term. To overcome these limitations, we introduce DiSCo (Distillation of Sparse Conversational retrieval), a novel approach that unifies retrieval and context modeling through a relaxed distillation objective. Instead of relying exclusively on representation learning, our method distills similarity scores between conversations and documents, providing more freedom in the representation space and better leveraging the contrastive nature of document relevance. Extensive experiments on Learned Sparse Retrieval (LSR) across five CS datasets demonstrate that DiSCo achieves substantial improvements in both in-domain and out-of-domain retrieval tasks, achieving up to a six-point gain in recall for out-of-domain datasets over state-of-the-art methods. Additionally, DiSCo employs a multi-teacher distillation strategy, using multiple LLMs as teachers, further enhancing performance and surpassing the individual teachers in in-domain settings. Furthermore, analysis of model sparsity reveals that DiSCo allows for more effective control over the sparsity of the trained models.
@@ 1. Introduction
Training retrieval models for this task is, however, challenging, because of the lack of large-scale conversational datasets, making human annotation an essential component of the process. In particular, a lot of effort has been invested first to create conversations with passage relevance judgment (Dalton et al., 2020b; Aliannejadi et al., 2024), and then to rewrite for each user utterance the contextualized version of each query (rewritten query) as the optimal denoised query of each turn (Elgohary et al., 2019). Datasets with rewrites enable us to learn the Conversational Query Rewriting (CQR) task by auto-regressive models (e.g., T5 (Raffel et al., 2019)), and then only pass the rewritten queries to retrieval models instead of the full noisy conversations. However, this two-step approach – rewrite and retrieve – is not efficient (Yu et al., 2021), and may lead to information loss and error within the rewrite phase that can propagate to the retrieval phase. Hence, the need for unified retrieval models that do both tasks together in the representation space (Penha and Hauff, 2020; Mo et al., 2024c).
@@
Approaches such as ConvDR (Yu et al., 2021), coSPLADE (Hai Le et al., 2023) and LeCoRe (Mao et al., 2023c) all learn both conversational context modeling and retrieval tasks within the representation space, either in a dense or sparse embedding space. As illustrated in Figure 1, they use a distillation objective enforcing the conversation representations to converge to the representations of gold human-rewritten queries (red arrow in the figure). This objective is, however, restrictive and assumes gold-rewritten query representations are the only optimal rewriting and only learning target in the representations space, leaving little freedom for the model to further learn and optimize the representations. DiSCo, differently from ConvDR relaxes the distillation targeting an entire hyperplane (blue hyperplane in the figure) instead of a single representation. This is achieved by focusing on query document similarity scores, rather than solely on the rewrite. Besides, previous work (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c) distills query rewrite independently from the ranking objective. We thus propose to fill this gap, with a single distillation loss that unifies both context modeling and ranking tasks, while not limiting the model to learn one single representation.
@@
Our method learns to distill the similarity scores between rewritten queries and documents rather than query representations directly. By distilling similarities, we first align with the contrastive nature of the ranking task (Tao and Zhai, 2007; Liu et al., 2021; Wang and Isola, 2020), but we also relax the training objective toward the final goal of retrieval, which is to compute similarities between queries and documents (Weinberger and Saul, 2009; Koch, 2015; Song et al., 2015). This relaxation allows the student to further learn the context modeling task, considering relevant and irrelevant documents from the corpus, rather than mapping all representations to the human representations. This objective also follows the precepts from Hofstätter et al. (2020) on distillation for ad-hoc retrieval and can benefit from hard negatives within the distillation loss. Both dense and sparse methods could be subject to this relaxation; however, in this work, we focus on learned sparse architectures (Zamani et al., 2018), such as SPLADE (Formal et al., 2021a; Lassance et al., 2024b), as the relaxation would have more degrees of freedom due to the high dimensionality of the representations for LSR.
@@
Besides, the proposed relaxation also reduces the level of constraints on the teachers, as any method producing similarity scores could be used for the distillation. In the context of CS, this allows the model to learn from multiple LLM rewrites, by fusing similarity scores of several teachers into a single score to be distilled. To the best of our knowledge, our work is the first to distill the knowledge of multiple LLM teachers for query rewriting. This also distinguishes our method from the work from Hofstätter et al. (2020), as we distill from LLM knowledge through the rewrites.
@@
We propose DiSCo, a relaxation of the training objective of CS model distilling similarities of rewritten queries rather than representations. 111Models checkpoints and code at https://github.com/SimonLupart/disco-conv-splade
We propose to distill knowledge from multiple teachers, unlocking the potential of mixtures of LLMs in CS.
We evaluate the effectiveness of DiSCo on in-domain (QReCC, TopiOCQA) and out-of-domain (CAsT 2020, 2022, and iKAT 2023) datasets, achieving state-of-the-art performance.
@@
Our results demonstrate that our Distillation of Sparse Con-versational retrieval, DiSCo, leads to in-domain performance improvements, through our relaxation of the distillation loss. We also see improved generalization capacities with 12% gains on recall and 16% on precision compared to previous zero-shot models (LeCoRe and QRACDR) on CAsT 2020. We demonstrate that distilling from multiple LLM teachers for the query rewriting task brings further performance gains, compared to single LLM teacher. Finally, thanks to the increased freedom on the representations, we show that we can better control the model sparsity, even for long contexts, in deeper turns of the conversations.
@@ 2. Related Work
Conversational Search (CS) differentiates itself from ad-hoc search primarily through the nature of the input. Conversational history can grow very long with turns’ dependencies, whereas search engine queries are typically concise, limited to a few words (Cambazoglu et al., 2021; Lupart et al., 2023). The main challenge in addressing CS is modeling the conversational context, to only represent useful information from the previous turns (Wu et al., 2021). Two possibilities exist to learn this noise reduction: first on the token level, with generative models that learn to generate contextualized queries from past conversations (Elgohary et al., 2019), or within the representation spaces (Yu et al., 2021) of retrieval models (Mo et al., 2024b).
@@
Query rewriting. Learning to resolve the conversational history is difficult, as no or very few conversational search engines are in production today, limiting the availability of user-interaction data. The effort is thus mostly based on human-annotated data. One of the first conversational datasets are QuAC (Choi et al., 2018) and ORConvQA (Qu et al., 2020), where human annotators were tasked to create conversations out of existing documents, assuming a human-machine conversation over the topic of the document. Each paragraph could potentially be an utterance and relevant passages would be inferred by construction. Later CANARD (Elgohary et al., 2019) was released on top of ORConvQA, consisting of human rewrites of each utterance of the conversations, to help learn the context modeling task. This defined one of the main sub-tasks of CS: Conversational Query Rewriting (CQR) (Elgohary et al., 2019; Ye et al., 2023). Following studies learned the CQR task at the token level – generating automatic query rewrites based on the conversation history – on autoregressive language models (e.g., T5 or BART) (Mao et al., 2023a; Wu et al., 2021). Similar methods (Mo et al., 2023) used language models to answer the query, before resolving the context. Today, work is still being done to solve the CQR task, using LLMs in zero- or few-shot fashion. CHIQ (Mo et al., 2024a) proposes to decompose the context modeling task into several simpler sub-tasks for the LLMs – history enhancement, answer generation, question disambiguation – to gain in interpretability and effectiveness, LLM4CS aggregates several rewrites and answers within the representation space of the retrieval models (Mao et al., 2023b), and MQ4CS focuses on multi-aspect query rewrites (Abbasiantaeb et al., 2024). However, context modeling using LLMs is too computationally costly, making it impossible for production.
@@
Distillation in CS. Lots of research aims at modeling the conversational context via learning to represent the conversation based on gold rewritten queries. Existing methods achieve this goal by taking the representation of the rewritten gold query as a teacher model and learning to distill the representation (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024c). ConvDR (Yu et al., 2021) learns the mapping between full conversation and human rewrites representations by minimizing an Mean Square Error (MSE) loss on the CLS tokens of both, performing well for dense bi-encoders. QRACDR (Mo et al., 2024c), also proposes a similar distillation, with several new MSE terms, between documents and queries to improve the representations and better align with the contrastive nature of the task. We differentiate from these works as we inspire from contrastive margin distillation (Margin-MSE (Hofstätter et al., 2020)) and relax the query rewriting distillation by learning based on the relevance scores (i.e., the dot product of the query representation and a document), rather than the representation itself. This way, not only do we enable converging towards indefinite possible query representations, but also we take advantage of the contrastive nature of ranking by aligning the rewrite task to retrieval. Furthermore, we are not bound to one query or one teacher.
@@
Learned sparse representation. As Learned Sparse Retrieval (LSR) gained popularity in ad-hoc retrieval (Lassance et al., 2024b; Lassance and Clinchant, 2023, 2022b), SPLADE architectures were proposed for conversational passage retrieval (Hai Le et al., 2023; Mao et al., 2023c; Lupart et al., 2024), to benefit their interpretability and robustness properties (Formal et al., 2021b, 2022; Lupart and Clinchant, 2023). In coSPLADE (Hai Le et al., 2023), the authors use an MSE loss between full conversation and human rewrite representations of the sparse bag-of-words representations (of dimension 30k), and show promising performance. In the meantime, LeCoRe (Mao et al., 2023c) also aims to distill human rewrites on sparse representations, through intermediate embedding layers and by filtering a maximum of dimensions, to avoid the MSE on large dimensionality vectors. Distilling sparse representations is, however, challenging, as it requires controlling which dimensions should be activated, often restricting the potential of the sparse representations. Additionally, using a MSE is problematic since it treats each dimension separately, neglecting the significance of the associated dimension. These issues limit the effectiveness of sparse retrieval in previous work. In our work, by only distilling end scores, we give the model complete freedom to learn which dimensions to activate. We also investigate sparsity further, including methods to control it within CS.
@@ 3.1. Preliminaries
Notation. We consider a set of conversations between a user and a system, each composed of multiple turns. At turn , we have access to previous queries and answers, together with the final user utterance . We denote the full conversational context as , separated with [SEP] tokens, and as the gold rewritten query. resolves various complexities such as ellipsis and ambiguities of the conversation and can be generated by either human or LLM.
Furthermore, similarly to other approaches in CS (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024c) we rely on two already trained encoder models and , for queries and documents representations. These backbone models are trained on a large-scale IR dataset with regular short-form queries extracted from search engine logs. We also follow the assumption made in CS that is already good at representing documents and doesn’t need further fine-tuning, only needs further fine-tuning to adapt to long conversational contexts, noted . This can be done with a contrastive InfoNCE (Oord et al., 2018) loss on the CS datasets (e.g., convANCE, convSPLADE), or with a distillation loss (e.g., ConvDR and other distillation approaches such as DiSCo).
Learned Sparse Retrieval. We rely on the SPLADE architecture (Formal et al., 2021b, a), based on the BERT pretrained transformers (Devlin et al., 2018; Lin et al., 2020). SPLADE makes use of the Masked Language Modeling embedding layer of BERT to create sparse representations, where each dimension corresponds to a token of the vocabulary. , and are the MLM outputs of the model, of dimension 30k (vocabulary dimension). To control the sparsity of the model, the authors propose to use a regularization loss (Paria et al., 2020), The sparsity of the models can be controlled with the two hyperparameters and , as weights for the regularization loss. As we rely on the same SPLADE model, we also included this loss.
@@ 3.2. DiSCo
Representation distillation. Existing approaches (Yu et al., 2021; Hai Le et al., 2023; Mao et al., 2023c; Mo et al., 2024c) distill query rewrites on the representation space, with a direct constraint on the representation. The goal is to have the representations of the full conversations converging toward the representations of the rewritten queries:
@@
with targeted scores , and the dimension of the representation space. More intuitively, while the existing distillation loss functions bound the model to converge to one single embedding (Yu et al., 2021), our loss allows for infinite possible optimum embeddings, as long as they have the same dot product with the target relevant document embedding (i.e., any point in the hyperplane). Such distillation on scores can be trained with a Kullback Leibler Divergence loss (Kullback and Leibler, 1951) on the distribution of scores:
@@
Hard negatives. As our new training objective from Equation 3 is now contrastive, we can benefit from hard negative mining. This was not possible with the previous distillation objective, as it only uses conversation representations, independently from documents.
Thus applied to the case of multiple negatives per query, the target space from Equation 2 becomes a subspace of (h-N) dimensions. This still keeps a higher degree of freedom, since the number of negatives is much lower than the dimension of the embedding space ().
@@
Teacher Models. An important component of distillation relies on the choice of the teacher. The teacher is first used to rewrite the conversation utterance, into , which will be encoded and distilled through the similarities scores . We use multiple LLMs, together with human rewrite (when available), as teachers. Having several teachers is motivated by the well-established effectiveness of ensembling strategies, where stronger teachers would provide higher-quality training signals, leading to better student model performance. This results in several trained DiSCo students, depending on the teacher used. Below we list all the teachers:
T0: T5QR
@@
Furthermore, another property of the proposed distillation is that we can distill from multiple teachers, making the distill from several LLMs or humans possible. We do so by averaging the similarity scores of the teachers, each having a different rewrite, thus a different similarity with the documents. We experimented with mean, min, and max aggregation methods, observing no significant differences in performance, and thus decided to use the more simple mean aggregation. Overall, the distilled score is:
@@
Student-Teacher Fusion. Finally, we propose a fusion of the teacher and the student model, combining the ranked lists of both at inference. This is achieved using an average normalized score over the ranked list of both models.